# Analytical and Numerical Studies of Dark Current in Radiofrequency Structures for Short-Pulse High-Gradient Acceleration

### v2

@@ I Background and Introduction
Achieving a high accelerating gradient is crucial for compact linear accelerators, but the presence of radiofrequency (RF) breakdown poses a significant challenge. Recently, acceleration with RF pulses on the order of a few nanoseconds has emerged as a promising approach for overcoming long-standing limitations in the gradient achievable by mitigating the impact of RF breakdown [1, 2, 3, 4, 5, 6, 7, 8]. These short RF pulses have been generated by various approaches including RF pulse compressors [9, 10, 11], and power extractors [12, 13, 14, 15, 16] based on the structure wakefield acceleration (SWFA) concept [17]. SWFA is one promising advanced accelerator concept, where a high-charge drive beam excites wakefield in metallic or dielectric structures in vacuum, and then the excited wakefield can be used to accelerate a main beam at high gradients. There are two schemes of SWFA: collinear wakefield acceleration, where the drive and main beams travel in the same structure, whereas in two-beam acceleration (TBA) the beamlines are decoupled. In the TBA scheme, a power extractor is designed to extract the wakefield generated by the drive beam, often a train of high-charge electron bunches. Each electron bunch in the train excites a short RF pulse as it traverses the power extractor, and when the wakefield pulses generated from multiple electron bunches are coherently added, a combined RF pulse with high peak power can be produced. Using this approach, RF pulses with a duration of a few nanoseconds and a peak power of over 500 MW have been generated at $X$-band [14, 15]. These short RF pulses can then be transferred and coupled into accelerating structures, where gradients exceeding the limit in conventional accelerators (100 MV/m for $X$-band for example) have been demonstrated in a series of experiments [18].
@@
One of these experiments at the Argonne Wakefield Accelerator (AWA) facility successfully demonstrated an electric field approaching 400 MV/m on the photocathode surface in an $X$-band photogun with (1+1/2)-cell resonant cavities [5]. These cavities are powered by approximately 9 ns long RF pulses extracted from an eight-electron-bunch train with a total charge of about 400 nC using an $X$-band power extractor [15]. Figure 1(a) shows the photogun cavities, comprising a full cell and a half cell, and an input waveguide port. A coaxial coupler was designed to maintain good field symmetry at the photocathode. A directional coupler (not shown) was attached to the waveguide port to measure the input and reflected RF signals. The cavities operate in the TM010,π mode with a center frequency of 11.7 GHz, with the electric field distribution shown in Fig. 1(b). The input RF signal exhibits a time profile of approximately 3 ns rise, 3 ns flattop and 3 ns fall, as shown in Fig. 1(c). This time profile is determined by the design of the power extractor and the drive bunch train length. The electric field at the center of the photocathode for an input pulse with a peak power of 200 MW is also shown in Fig. 1(c).
@@
Figure 1: Model and RF design of the $X$-band photogun cavities driven by short RF pulses. (a) Cross-sectional view of the copper cavities along the central plane. (b) Normalized electric field distribution at 11.7 GHz. (c) Electric field on the photocathode surface for an input RF pulse with 200 MW peak power; the inset shows the time profile of the input RF signal, with a 3 ns rise, 3 ns flattop and 3 ns fall.
@@
The encouraging high gradient demonstrated experimentally points to the need for a deeper physics understanding of RF breakdown processes in the under-explored parameter space of $\mathcal{O}\!\bigl{(}1\,\text{ns}\bigr{)}$ pulses. An empirical scaling law [19], established from breakdown studies in klystron-powered accelerating structures with $\mathcal{O}\!\bigl{(}100\,\text{ns}\bigr{)}$ RF pulse durations, relates the breakdown rate (BDR) to the accelerating gradient $E_{a}$ and the pulse length $t_{p}$ as $E_{a}^{30}\times t_{p}^{5}/\text{BDR}=\text{constant}$, suggesting that higher gradients can be achieved with shorter pulses. However, recent experiments using $\mathcal{O}\!\bigl{(}1\,\text{ns}\bigr{)}$ long pulses revealed a different scaling behavior [7, 6] and interesting dark current behavior [5]. During the RF conditioning of the short-pulse $X$-band photogun, very low dark current was observed. Despite monitoring with both a Faraday cup and an integrating current transformer (ICT), no clear signal was detected, suggesting an upper limit of approximately 1 pC in dark charge. Dark-current-induced beam loading was also observed in this experiment, as shown in Fig. 2(a). The ratio $R$ of the measured and simulated peak reflected RF signal was close to unity in the absence of RF breakdown or dark current [Figs. 2(b) and 2(d)], but dropped below unity during a stage where beam loading occurred [Fig. 2(c)]. These observations call for further investigation of the dark current dynamics in the short-pulse regime.
@@
Dark current refers to the unwanted flow of electrons in the absence of an injected beam [20], which can coexist with the primary beam and cause instabilities, power losses, and structural damage due to electron bombardment [21]. The growth of dark current may also contribute to RF breakdown, which typically manifests as vacuum arcs that can disrupt RF power flow and may even cause irreversible damage to accelerating structures. RF breakdown is strongly influenced by the localized electromagnetic field on the inner surfaces of the structure, as well as by interactions between these fields and dark electrons [22]. A range of physical processes across various timescales are associated with dark current dynamics.
@@
Figure 2: RF conditioning of the $X$-band photogun cavities. (a) Conditioning history of the peak electric field $E_{0}$ at the photocathode surface (blue), and the ratio of the measured to simulated peak reflected RF signal $R$ (orange). (b-d) Reflected RF signals at the corresponding stages denoted by the vertical lines in Panel (a). This figure is reprinted with permission from Tan et al., Phys. Rev. Accel. Beams 25, 083402 (2022) under the terms of the Creative Commons Attribution 4.0 International license.
@@
A common source of dark electrons is field emission from surface regions with high local fields [23, 24]. The local field $E_{\text{local}}$ is often significantly higher than the applied macroscopic field $E_{\text{applied}}$ due to surface irregularities, and can be expressed as $E_{\text{local}}=\beta E_{\text{applied}}$, where $\beta$ is the field enhancement factor. Material properties such as work function, roughness, and composition can strongly influence field emission [25, 26, 27, 28]. Field emission commonly occurs in high-field regions, where pulsed heating – localized surface temperature rise via Joule heating – can vaporize emitting sites, triggering explosive electron emission and surface modification, which may in turn further enhance field emission [29]. Electron multipacting is another important mechanism contributing to dark current growth [30, 31, 32, 33, 34, 35, 36, 37, 38]. In this process, emitted electrons from the surface interact with the RF field and may return to the surface again, releasing secondary electrons. Under certain resonant conditions, this process repeats, leading to exponential electron multiplication. Plasma can form as the electron density increases [39, 40, 41]. This plasma can short-circuit RF structures, trigger further ionization, and escalate into arcing.
@@
We report analytical and numerical studies on the dynamics of dark current in the short-pulse acceleration regime, using the $X$-band cavities shown in Fig. 1 as a representative case. Section II presents an analytical theory of electron multipactor, where we identify the multipacting resonance conditions from electron trajectory calculations and evaluate the secondary electron yield (SEY). Section III describes simulations of field emission and multipacting processes using the cst [42] Particle Studio, analyzing the impact of pulse shape, pulse length and structure gradient on dark current growth. Section IV discusses the formation of multipacting electron clouds and their influence on RF pulse distortion. Finally, Sec. V summarizes the key findings and outlines future directions for high-gradient short-pulse RF acceleration research.
@@ II Theory of Electron Multipacting
Previous analytical models have explored electron multipacting behavior in various RF configurations [43, 44, 45, 46, 47, 48]. For example, Ref. [43] presents a theory of single-sided multipacting on a dielectric surface with an RF electric field and a DC electric field perpendicular to the dielectric surface. The DC electric field is necessary to extract electrons from the dielectric. Ref. [44] investigates the effect of the RF magnetic field on multipacting on dielectric surfaces. These results are important for evaluating multipacting behavior in dielectric windows. Multipacting on metallic surfaces of accelerator structures has also been studied. Ref. [45] investigates two-sided multipacting between two metal electrodes. Ref. [46] analyzes the single-sided multipacting on cavity sidewalls under the influence of normal electric fields. The study identifies electron trajectories that return to the emission surface and contribute to sustaining the discharge.
In this work, we study electron multipacting in crossed RF fields near the sidewall of a pillbox cavity, focusing on closed electron trajectories that originate from and return to the sidewall. These trajectories are of particular interest, as they are resonant with the RF field and are expected to play a dominant role in the multipactor phenomenon. Our analysis includes an axial electric field $E_{z}$, a radial electric field $E_{r}$, and an azimuthal magnetic field $B_{\theta}$. The radial component $E_{r}$ arises near the cavity sidewall due to the presence of the cavity iris. We modify the field distribution in the pillbox cavity to approximate conditions near the sidewall region of the $X$-band photogun cavities. We assume an axisymmetric field distribution, so electron motion is confined to the rrzz plane. This assumption holds as long as any asymmetries or perturbations that could induce out-of-plane motion remain negligible. These approximations yield predictions in good agreement with simulations of the actual cavity, as presented in Sec. III. We begin by calculating secondary electron trajectories and then evaluate the SEY in both single-impact and multipacting resonance conditions.
Figure 3: Schematic representation of the multipactor analysis setup. (a) Simulation geometry used for electron trajectory evaluation. (b) Crossed RF field configuration showing an electron trajectory under combined $E_{r}$, $E_{z}$, and $B_{\theta}$ fields. The parameters $v_{0}$ (initial velocity) and $\alpha$ (emission angle) are defined at emission, while $r$ and $z$ denote the radial and longitudinal coordinates, respectively; $a$ and $L$ are the cavity radius and length, respectively.
@@ II.1 Secondary electron trajectory calculations
Under the assumptions outlined above, we investigate electron multipacting near the sidewall of a pillbox cavity. The cavity has a radius of $a=9.49$ mm and a length of $L=6.50$ mm, corresponding to the full-cell dimensions of the $X$-band photocathode cavities. The configuration used for the electron trajectory calculations is illustrated in Fig. 3. The origin, $r=z=0$, is defined at the center of the pillbox cavity. An electron is emitted from the sidewall at $t=0$ with an emission angle $\alpha_{0}$, initial velocity $v_{0}$, and initial RF phase $\phi_{0}$. The initial axial position is $z(t=0)=z_{0}$, and the radial distance from the sidewall is $\Delta r(t=0)=r(t=0)-a=0$. We assume a steady-state field, with the axial electric field amplitude at the cavity center denoted by $E_{z0}$. The equations of motion for the electron are:
@@
[…]
@@
$E_{r}(\Delta r,z)=E_{r_{0}}\left(1+S\cdot\Delta r\right)\cos\left(\frac{\pi z}{L}\right)\cos(\omega t+\phi_{0}),$ (3)
where $E_{r0}$ is the peak amplitude of $E_{r}$, $\omega$ is the angular frequency of the RF field, and $S=dE_{r}/dr$, fitted from the simulated field distribution. The axial electric field $E_{z}$ is:
@@
$B_{\theta}(\Delta r,z)=\left[(B_{0}+S_{1}\cdot\Delta r)+(B_{1}+S_{2}\cdot\Delta r)\right.\\ \left.\times\sin\left(\frac{\pi z}{L}+\zeta\right)\right]\sin(\omega t+\phi_{0}),$ (5)
@@
While we employ the above field distributions to calculate secondary electron trajectories for the remainder of this section, we have also developed a simplified theory neglecting the radial variation of $E_{z}$ and $B_{\theta}$. This approximation yields closed-form expressions for multipacting resonance conditions, useful for rapid design estimates. The simplified theory, presented in Appendix A, agrees well with the full model in the low-field regime. For accurate trajectory tracking and SEY evaluation, however, we retain full radial dependence in all RF field components.
II.2 Single-cycle SEY calculations
The SEY, $\delta$, quantifies the number of secondary electrons emitted per incident electron. Here, we evaluate $\delta$ after a single emission–return cycle, defined as the electron’s first re-impact with the cavity sidewall following its initial emission.
@@ II.2 Single-cycle SEY calculations
Figure 5: Single-cycle SEY ($\delta$) map for the full cell of the X-band photogun cavities for different combinations of initial axial emission position $z_{0}$ and RF field amplitudes, characterized by the axial electric field at the cavity center $E_{z0}$. The color scale indicates the magnitude of $\delta$.
[…]
Figure 6: Sample multipacting resonance modes ($N=1\text{--}7$) calculated for the RF field distribution in the full cell of the $X$-band photocathode cavities. The corresponding axial field amplitudes $E_{z0}$ at the cavity center for each order are 106, 61, 39, 31, 26, 22, and 19 MV/m, respectively.

### v3

@@ I Background and Introduction
Achieving a high accelerating gradient is crucial for compact linear accelerators, but the presence of radiofrequency (RF) breakdown poses a significant challenge. Recently, acceleration with RF pulses on the order of a few nanoseconds has emerged as a promising approach for overcoming long-standing limitations in the gradient achievable by mitigating the impact of RF breakdown [1, 2, 3, 4, 5, 6, 7, 8]. These short RF pulses have been generated by various approaches including RF pulse compressors [9, 10, 11], and power extractors [12, 13, 14, 15, 16] based on the structure wakefield acceleration (SWFA) concept [17]. SWFA is one promising advanced accelerator concept, where a high-charge drive beam excites wakefield in metallic or dielectric structures in vacuum, and then the excited wakefield can be used to accelerate a main beam at high gradients. There are two schemes of SWFA: collinear wakefield acceleration, where the drive and main beams travel in the same structure, and two-beam acceleration (TBA), where the two beamlines are decoupled. In the TBA scheme, a power extractor is designed to extract the wakefield generated by the drive beam, often a train of high-charge electron bunches. Each electron bunch in the train excites a short RF pulse upon traversing the power extractor, and when the wakefield pulses generated from multiple electron bunches are coherently added, a combined RF pulse with high peak power can be produced. Using this approach, RF pulses with a duration of a few nanoseconds and a peak power of over 500 MW have been generated at $X$-band [14, 15]. These short RF pulses can then be transferred and coupled into accelerating structures, where gradients exceeding the limit in conventional accelerators (100 MV/m for $X$-band for example) have been demonstrated in a series of experiments [18].
@@
One of these experiments at the Argonne Wakefield Accelerator (AWA) facility successfully demonstrated an electric field approaching 400 MV/m on the photocathode surface in an $X$-band photogun with (1+1/2)-cell resonant cavities [5]. These cavities are powered by approximately 9 ns long RF pulses extracted from an eight-electron-bunch train with a total charge of about 400 nC using an $X$-band power extractor [15]. Figure 1(a) shows the photogun cavities, comprising a full cell and a half cell, and an input waveguide port. A coaxial coupler was designed to maintain good field symmetry at the photocathode. A directional coupler (not shown) was attached to the waveguide port to measure the input and reflected RF signals. The cavities operate in the TM010,π mode with a center frequency of 11.7 GHz, and the electric field distribution is shown in Fig. 1(b). The input RF signal exhibits a time profile of approximately 3 ns rise, 3 ns flat top and 3 ns fall, as shown in Fig. 1(c). This time profile is determined by the design of the power extractor and by the spacing and total length of the drive bunch train. The electric field at the center of the photocathode for an input pulse with a peak power of 200 MW is also shown in Fig. 1(c).
@@
Figure 1: Model and RF design of the $X$-band photogun cavities driven by short RF pulses. (a) Cross-sectional view of the copper cavities along the central plane. (b) Normalized electric field magnitude distribution at 11.7 GHz. (c) Electric field on the photocathode surface for an input RF pulse with 200 MW peak power; the inset shows the time profile of the input RF signal, with a 3 ns rise, 3 ns flat top and 3 ns fall.
@@
The encouraging high gradient demonstrated experimentally points to the need for a deeper physics understanding of RF breakdown processes in the under-explored parameter space of $\mathcal{O}\!\bigl(1\,\text{ns}\bigr)$ pulses. An empirical scaling law [19], established from breakdown studies in klystron-powered accelerating structures with $\mathcal{O}\!\bigl(100\,\text{ns}\bigr)$ RF pulse durations, relates the breakdown rate (BDR) to the accelerating gradient $E_{a}$ and the pulse length $t_{p}$ as $E_{a}^{30}\times t_{p}^{5}/\text{BDR}=\text{constant}$, suggesting that higher gradients can be achieved with shorter pulses. However, recent experiments using $\mathcal{O}\!\bigl(1\,\text{ns}\bigr)$ long pulses revealed different scaling [7, 6] and distinct dark current signatures [5]. During the RF conditioning of the short-pulse $X$-band photogun, very low dark current was observed. Despite monitoring with both a Faraday cup and an integrating current transformer, no clear signal was detected, suggesting an upper limit of approximately 1 pC on the dark-current charge at the location of the monitors [5]. Dark-current-induced beam loading was also observed in this experiment, as shown in Fig. 2(a). The ratio $R$ of the measured and simulated peak reflected RF signals was close to unity in the absence of RF breakdown or dark current [Figs. 2(b) and 2(d)], but dropped below unity during a stage where beam loading occurred [Fig. 2(c)]. These observations call for further investigation of the dark current dynamics in the short-pulse regime.
@@
Dark current refers to the unwanted flow of electrons in the absence of an injected beam [20], and it can coexist with the primary beam and cause instabilities, power losses, and structural damage due to electron bombardment [21]. The growth of dark current may also contribute to RF breakdown, which typically manifests as vacuum arcs that can disrupt RF power flow and may even cause irreversible damage to accelerating structures. RF breakdown is strongly influenced by the localized electromagnetic field on the inner surfaces of the structure, as well as by interactions between these fields and dark electrons [22]. A range of physical processes across various timescales are associated with dark current dynamics.
@@
Figure 2: RF conditioning of the $X$-band photogun cavities. (a) Conditioning history of the peak electric field $E_{0}$ at the photocathode surface (blue), and the ratio of the measured to simulated peak reflected RF signal $R$ (orange). (b-d) Comparison of measured (dashed red) reflected RF signals with simulated signals (solid green, assuming no RF breakdown) at the corresponding stages denoted by the vertical lines in Panel (a). This figure is reprinted with permission from Tan et al., Phys. Rev. Accel. Beams 25, 083402 (2022) under the terms of the Creative Commons Attribution 4.0 International license.
@@
A common source of dark electrons is field emission from surface regions with high local fields [23, 24]. The local field $E_{\text{local}}$ is often significantly higher than the applied macroscopic field $E_{\text{applied}}$ due to surface irregularities, and can be expressed as $E_{\text{local}}=\beta E_{\text{applied}}$, where $\beta$ is the field enhancement factor. Material properties such as work function, roughness, and composition can strongly influence field emission [25, 26, 27, 28]. Field emission commonly occurs in high-field regions, where pulsed heating – localized surface temperature rise via Joule heating – can vaporize emitting sites, triggering explosive electron emission and surface modification, which may in turn further enhance field emission [29]. Electron multipacting is another important mechanism contributing to dark current growth [30, 31, 32, 33, 34, 35, 36, 37, 38]. In this process, emitted electrons from the surface interact with the RF field and may return to the surface again, releasing secondary electrons. Under certain resonant conditions, this process becomes self-sustaining and can lead to exponential electron multiplication. As the electron density increases, a plasma may form [39, 40, 41].The resulting plasma can short-circuit RF structures, trigger further ionization, and escalate into arcing.
@@
We report analytical and numerical studies on the dynamics of dark current in the short-pulse acceleration regime, using the $X$-band cavities shown in Fig. 1 as a representative case. Section II presents an analytical theory of electron multipactor, where we identify the multipacting resonance conditions from electron trajectory calculations and evaluate the secondary electron yield (SEY). Section III describes simulations of field emission and multipacting processes using the cst [42] Particle Studio, analyzing the impact of pulse shape, pulse length, and structure gradient on dark current growth, with a focus on comparing the long- and short-pulse regimes. Section IV discusses the formation of multipacting electron clouds and their influence on RF pulse distortion. Finally, Sec. V summarizes the key findings and outlines future directions for high-gradient short-pulse RF acceleration research.
@@ II Theory of Electron Multipacting
Previous analytical models have explored electron multipacting in various RF configurations [43, 44, 45, 46, 47, 48]. For example, Ref. [43] presents a theory of single-sided multipacting on a dielectric surface with an RF electric field parallel to the dielectric surface and a DC electric field perpendicular to the surface. Ref. [44] investigates the effect of the RF magnetic field on multipacting on dielectric surfaces. These results are important for evaluating multipacting behavior in dielectric windows. Multipacting on metallic surfaces of accelerator structures has also been studied. Ref. [45] investigates two-sided multipacting between two metal electrodes. Ref. [46] analyzes the single-sided multipacting on cavity sidewalls under the influence of crossed fields. The study identifies electron trajectories that return to the emission surface and contribute to sustained discharge.
In this work, we investigate electron multipacting in crossed RF fields near the sidewall of a pillbox cavity, focusing on closed electron trajectories that originate from and return to the sidewall. These trajectories are of particular interest, as they are resonant with the RF field and are expected to play a dominant role in the multipacting process. Our analysis includes an axial electric field $E_{z}$, a radial electric field $E_{r}$, and an azimuthal magnetic field $B_{\theta}$. The radial component $E_{r}$ arises near the cavity sidewall due to the presence of the cavity iris. We modify the field distribution in the pillbox cavity to approximate conditions near the sidewall region of the $X$-band photogun cavities. We assume an axisymmetric field distribution, so electron motion is confined to the rrzz plane. This assumption holds as long as any asymmetries or perturbations that could induce out-of-plane motion remain negligible. These approximations yield predictions in good agreement with simulations of the realistic cavity, as presented in Sec. III.2. We begin by calculating secondary electron trajectories and then evaluate the SEY under both single-trip and multipacting resonance conditions.
Figure 3: Schematic representation of the multipactor analysis setup. (a) Simulation geometry used for electron trajectory evaluation. (b) Crossed RF field configuration showing an electron trajectory under combined $E_{r}$, $E_{z}$, and $B_{\theta}$ fields. The parameters $v_{0}$ (initial velocity) and $\alpha$ (emission angle) are defined at emission. $r$ and $z$ denote the radial and longitudinal coordinates, respectively, while $a$ and $L$ are the cavity radius and length, respectively.
@@ II.1 Secondary electron trajectory calculations
Under the assumptions outlined above, we investigate electron multipacting near the sidewall of a pillbox cavity. The cavity has a radius of $a=9.49$ mm and a length of $L=6.50$ mm, corresponding to the full-cell dimensions of the $X$-band photocathode cavities. The configuration used for the electron trajectory calculations is illustrated in Fig. 3. The origin, $r=z=0$, is defined at the center of the pillbox cavity. An electron is emitted from the sidewall at $t=0$ with an emission angle $\alpha_{0}$, initial velocity $v_{0}$, and initial RF phase $\phi_{0}$. The initial axial position is $z(t=0)=z_{0}$, and the radial distance from the sidewall is $\Delta r(t=0)=r(t=0)-a=0$. We assume a steady-state RF field, with the axial electric field amplitude at the cavity center, $E_{z0}$, treated as constant. The equations of motion for the electron are:
@@
Figure 4: Comparison of RF field amplitudes from cst simulations and analytical approximations for the full-cell cavity with radius $a=9.49$ mm and length $L=6.50$ mm. The field amplitudes, $B_{\theta}$ in (a), $E_{r}$ in (b), and $E_{z}$ in (c), are all scaled to an axial field amplitude of $E_{z0}=100$ MV/m at the cavity center. Solid red curves represent cst simulation data, and black dashed lines indicate the analytical field amplitudes used in the model. The shaded region in (c) indicates the radial range near the sidewall where the analytical field approximation is applied.
@@
$E_{r}(\Delta r,z)=E_{r_{0}}\left(1+S\Delta r\right)\cos\left(\frac{\pi z}{L}\right)\cos(\omega t+\phi_{0}),$ (3)
where $E_{r0}$ is the peak amplitude of $E_{r}$, $\omega$ is the angular frequency of the RF field, and $S=(dE_{r}/dr)/E_{z0}$, fitted from the simulated field distribution. The axial electric field $E_{z}$ is:
@@
$B_{\theta}(\Delta r,z)=\left[(B_{0}+S_{1}\Delta r)+(B_{1}+S_{2}\Delta r)\right.\\ \left.\times\sin\left(\frac{\pi z}{L}+\zeta\right)\right]\sin(\omega t+\phi_{0}),$ (5)
@@
The above model with approximate field distributions uses a small set of fitted parameters and can be rapidly adapted to structures with similar field profiles. While we employ the above field distributions to calculate secondary electron trajectories for the remainder of this section, we have also developed a simplified theory neglecting the radial variation of $E_{z}$ and $B_{\theta}$. This approximation yields closed-form expressions for multipacting resonance conditions, useful for rapid design estimates. The simplified theory, presented in Appendix A, agrees well with the full model at low fields. For accurate trajectory tracking and SEY evaluation, however, we retain full radial dependence in all RF field components.
II.2 Single-trip SEY calculations
We define a single-trip multipacting trajectory as the initial excursion from emission to the first return impact on any surface, regardless of RF period count elapsed or whether the electron returns to its emission site. This contrasts with the multi-cycle resonances in Sec. II.3, where repeated resonant electron impacts occur over multiple RF cycles. Here we evaluate the SEY, $\delta$, as the number of secondary electrons emitted per incident electron, after a single emission–return trip.
@@ II.2 Single-trip SEY calculations
Figure 5: Single-trip SEY ($\delta$) map for the full cell of the $X$-band photogun cavities for different combinations of initial axial emission position $z_{0}$ and RF field amplitudes, characterized by the axial electric field at the cavity center $E_{z0}$. The color scale indicates the magnitude of $\delta$.
To compute the SEY, we begin by extracting the impact energy $W_{i}$ and incident angle $\theta_{i}$ from the electron trajectory. The axial impact location is denoted by $z_{i}$, and the RF phase at impact is given by $\phi_{i}=\omega\tau+\phi_{0}$, where $\tau$ is the time of flight. We assume an initial kinetic energy of 2 eV for secondary electrons, consistent with the emission model in Ref. [30]. We then perform a parameter sweep over the electron emission angle $\alpha_{0}$ and initial RF phase $\phi_{0}$, across a range of axial emission positions $z_{0}$ and field amplitudes characterized by the axial electric field amplitude at the cavity center, $E_{z0}$. All the field components ($E_{z}$, $E_{r}$, and $B_{\theta}$) are scaled proportionally with $E_{z0}$, which serves as the reference parameter for the sweep. SEY values exceeding unity are retained and averaged over both $\alpha_{0}$ and $\phi_{0}$ to generate a representative single-trip SEY contour map. The resulting map is shown in Fig. 5, highlighting regions in $E_{z0}$$z_{0}$ space that are susceptible to enhanced secondary emission and the potential onset of electron multipacting.
Figure 6: Sample multipacting resonance modes ($N=1\text{--}7$) calculated for the RF field distribution in the full cell of the $X$-band photocathode cavities. The corresponding axial field amplitudes $E_{z0}$ at the cavity center for each order are 106, 61, 39, 31, 26, 22, and 19 MV/m, respectively. Arrows indicate the direction of electron motion along the trajectories.